Solve the system using the elimination or substitution method. x + 2y - 3z = 9; -x + 3y - 2z = 1; 2x - 5y + 2z = 2
Answer by checkley79(3341)

x + 2y - 3z = 9
-x + 3y - 2z = 1
2x - 5y + 2z = 2 ADD
------------------
2X+0Y-3Z=12
2X-3Z=12
MULTIPLY THE FIRST EQUATION BY 3.
MULTIPLY THE SECOND EQUATION BY -2& ADD
3X+6Y-9Z=27
2X-6Y+4Z=-2
--------------
5X-5Z=25 MULTIPLY BY 2
2X-3Z=12 MULTIPLY BY -5 & ADD.
10X-10Z=50
-10X+15Z=-60
----------------
5Z=-10
Z=-10/5
Z=-2 ANS.
5X-5*-2=25
5X+10=25
5X=25-10
5X=15
X=15/5
X=3 ANS.
3+2Y-3*-2=9
3+2Y+6=9
2Y=9-3-6
2Y=0
Y=0 ANS.
PROOF:
2*3-5*0+2*-2=2
6+0-4=2
2=2